The length of the hypotenuse of a 30-60-90 triangle is 5. Find the perimeter.
Answer by stanbon(75887)

The length of the hypotenuse of a 30-60-90 triangle is 5. Find the perimeter.
------
Side opposite the 30 degree angle = (1/2)hypotenuse = 2.5
---
Side opposite the 60 degree angle = 2.5*sqrt(3) = 4.33
---
Perimeter = 5 + 2.5 + 4.33 = 11.83
===========================
Cheers,
Stan H.
================